OpenStudy (anonymous):

Which fraction in simplest form is equivalent to zero point four? four-sixteenths eighty thirty-seconds twelve-thirds twenty-four sixtieths

OpenStudy (tkhunny):

Very strange question. None of the options is in "simplest form". 0.4 = 2/5

OpenStudy (anonymous):

hmmm.

OpenStudy (tkhunny):

24/60 = 2/5 - but that doesn't clear up the "simplest form" problem.

OpenStudy (anonymous):

exactly.